Given the following measurements from John’s health assessment, answer the question given:

            End diastolic volume = 105 ml

            End systolic volume = 30 ml

            Heart rate = 70 bpm

When John exercises with maximum intensity, the highest heartbeat he gets is 150 bpm. What is his cardiac reserve?

        A). 5,250 ml          B). 11,250 ml    C). 7,350 ml      D). 6,000 ml

(0.1pts)

Question 80 - Given the following measurements from John’s health assessment, answer the question given:             End diastolic volume = 105 ml             End systolic volume = 30 ml             Heart rate = 70 bpm When John exercises with maximum intensity, the highest heartbeat he gets is 150 bpm. What is his cardiac reserve?         A). 5,250 ml          B). 11,250 ml    C). 7,350 ml      D). 6,000 ml

Select 5,250 ml as your answer

5,250 ml

Select 11,250 ml as your answer

11,250 ml

Select 7,350 ml as your answer

7,350 ml

Select 6,000 ml as your answer

6,000 ml

Select as your answer

None of the above
Answer #1

Cardiac reserve means the cardiac output in excess of the normal which can be achieved.

Cardiac output = Stroke volume * Heart rate

Stroke volume = 105-30 =75 ml

So normally heart rate's 70 bpm but can be increased to 150 bpm. Stroke volume remains the same.

So cardiac reserve = (150-70) * 75 = 6000 ml

So answer is D.